OpenStudy (anonymous):

Given the equation radical 8x+1 = 5, solve for x and identify if it is an extraneous solution. a. x = 1/4 , solution is not extraneous b. x =1/4 , solution is extraneous c. x = 3, solution is not extraneous d. x = 3, solution is extraneous

OpenStudy (yamyam70):

you mean \[\sqrt{8x +1}\] = 5 ?

OpenStudy (anonymous):

Is it the question ? \[sqrt{8x+1}\]=5 Please answer and get answer :)

OpenStudy (anonymous):

@yamyam70 yes

OpenStudy (yamyam70):

hello, if a linear equation has a radical , square both sides :)

OpenStudy (anonymous):

so it would be 64x+1=25

OpenStudy (yamyam70):

\[(\sqrt{8x + 1} )^2 = (5)^2\]

OpenStudy (yamyam70):

nope, by squaring both sides, take notice of the left hand side, we remove the radical.

OpenStudy (yamyam70):

\[\sqrt{8x + 1 } \times \sqrt{8x +1} = 8x +1 \]

OpenStudy (anonymous):

so the left side would stay 8x+1

OpenStudy (yamyam70):

aha!

OpenStudy (yamyam70):

so please continue, 8x + 1 = (5)^2 :)

OpenStudy (anonymous):

x=3?

OpenStudy (yamyam70):

How did you get that answer? :)

OpenStudy (anonymous):

subtract 1 from each side then divide by 8.

OpenStudy (anonymous):

what does extraneous solution mean?

OpenStudy (yamyam70):

Correct! How would you tell if it is extraneous or not?

OpenStudy (anonymous):

idk what that means.

OpenStudy (yamyam70):

We check the equation, by substituting the value of x. :)

OpenStudy (anonymous):

i got 5=5

OpenStudy (yamyam70):

if you check it and the results of the left hand side is equal to the right hand side it is not extraneous :) if it does not equal , it is extraneous :)

OpenStudy (anonymous):

oh ok thankyou!

OpenStudy (yamyam70):

Welcome to open study! please read the code of conduct if you have time :)

OpenStudy (yamyam70):

http://openstudy.com/code-of-conduct

OpenStudy (yamyam70):

you're welcome :)

OpenStudy (anonymous):

another question!

OpenStudy (yamyam70):

sure :)

OpenStudy (anonymous):

Solve for x, given the equation \[\sqrt{x-5}+7=11\] x = 21, solution is extraneous x = 21, solution is not extraneous x = 81, solution is extraneous x = 81, solution is not extraneous

OpenStudy (yamyam70):

I see then same procedure, square both sides :)

OpenStudy (anonymous):

x-5+49=121

OpenStudy (yamyam70):

oh wait, 7 isn't involved

OpenStudy (yamyam70):

one moment please

OpenStudy (anonymous):

ok

OpenStudy (yamyam70):

are you sure about the given choices?

OpenStudy (anonymous):

yes i got 77 but that answer was not there

OpenStudy (yamyam70):

I see, we got same, I'm trying a different approach, one moment

OpenStudy (anonymous):

ok

OpenStudy (yamyam70):

okay here goes,

OpenStudy (yamyam70):

\[\sqrt{x-5} +7 =11\]

OpenStudy (yamyam70):

we tranfer 7 to the right hand side :)

OpenStudy (yamyam70):

what would we have?

OpenStudy (anonymous):

\[\sqrt{x-5}=18\]

OpenStudy (anonymous):

and then do we square both sides?

OpenStudy (yamyam70):

wrong :) , when we tranfer 7 , we change the signs right? :)

OpenStudy (anonymous):

yes the right side is supposed to be 4

OpenStudy (yamyam70):

correct :) then we square both sides, the equation should look like \[\sqrt{x-5} = 4 \]

OpenStudy (anonymous):

i got x= 17

OpenStudy (yamyam70):

wrong :) \[(\sqrt{x-5?})^2 = (4)\]

OpenStudy (yamyam70):

(x−5?−−−−−√)2=(4)^2

OpenStudy (anonymous):

do i foil them?

OpenStudy (yamyam70):

wait nevermind that

OpenStudy (yamyam70):

\[(\sqrt{x-5})^2=(4)^2\]

OpenStudy (yamyam70):

^^^^^^ thats the correct one

OpenStudy (anonymous):

so how to i solve for that?

OpenStudy (yamyam70):

remember when we square , \[\sqrt{x-5}\] it will give us x-5 :)

OpenStudy (yamyam70):

\[(\sqrt{x-5}) (\sqrt{x-5}) = x-5\]

OpenStudy (anonymous):

x=21

OpenStudy (yamyam70):

correct :) , so check if it is extraneous or not :)

OpenStudy (anonymous):

it is not

OpenStudy (yamyam70):

Correct :)

OpenStudy (anonymous):

Given the equation \[\sqrt[-4]{x-3}=12\], solve for x and identify if it is an extraneous solution. x = 0, solution is not extraneous x = 0, solution is extraneous x = 12, solution is not extraneous x = 12, solution is extraneous

OpenStudy (yamyam70):

is that -4 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

would it be this \[x-3\frac{ 1 }{ -4}=12\]

OpenStudy (yamyam70):

hmm, I'm not quite sure about the rule here.

OpenStudy (yamyam70):

I would suggest too start another session. :)

OpenStudy (yamyam70):

I'm sorry , I can't help you with this one :)

OpenStudy (anonymous):

ok thanks for the help

OpenStudy (yamyam70):

Bye :)

OpenStudy (yamyam70):

Don't forget to close the question :)

OpenStudy (anonymous):

how do i do that?

OpenStudy (yamyam70):

On the left side, there is a timer, "bump in" beside it click close. :)